OpenStudy (anonymous):

Courtney reads around 110 words per minute. She procrastinated and now has less than a week to finish her reading assignment. If the book she needs to finish has another 820 pages left and each page has around 350 words, how many hours should she set aside to complete her assignment? Show your work to receive full credit.

OpenStudy (anonymous):

help please

OpenStudy (pawanyadav):

No. Of words = 820×350=287000 No . of days =7=7×24×60minutes=10080min. No. Of word she have to read per minute=287000÷10080min. Divide it by her reading speed you will get time in min. Then. Convert it to hour as per the question.

OpenStudy (anonymous):

thanks